OpenStudy (khally92):

please what is the derivative of p(t)= 500/1+e^-t. thanks

OpenStudy (cwrw238):

i'm finding this one a bit tricky - not at my best today i think you could do it using quotient rule: dp(t)/ dt = (1+e^-t)*0 - 500( -e^-t) ---------------------- (1 + e^-t)^2 = 500e^-t ------- (1 + e^-t)^2

OpenStudy (khally92):

well let me know what the derivative of e^-t is first?

OpenStudy (cwrw238):

using chain rule its -1 * e^-t = -e^-t

OpenStudy (khally92):

yeah sure thanks

OpenStudy (cwrw238):

ok - tell you what i'll check it out on wolfram alpha - that doesn't make mistakes